In 6 years, Sylvia will be 6/5 Hue's age. Four years ago, Sylvia was 10 years less than twice Hue's age. How old are Sylvia and Hue now?
Answer by lwsshak3(11628)

In 6 years, Sylvia will be 6/5 Hue's age. Four years ago, Sylvia was 10 years less than twice Hue's age. How old are Sylvia and Hue now?
**
let x=Hue's present age
let y=Sylvia's present age
..
In 6 years:
Hue's age=(x+6)
Sylvia's age=(y+6)
(y+6)=(6/5)(x+6)
y=(6/5)(x+6)-6
y=(6x+36)/5-6
..
4 years ago:
Hue's age=(x-4)
sylvia's age=(y-4)
y-4=2(x-4)-10
y=2x-8-10+4
y=2x-14
..
(6x+36)/5-6=2x-14
LCD+5
6x+36-30=10x-70
4x=76
x=19
y=2x-14=38-14=24
..
ans:
Hue's present age=19
Sylvia's present age=24